OpenStudy (abbles):

What is the binomial coefficent of the 5th term of (x + y)^10

OpenStudy (nincompoop):

how about you try solve it and show us what you know at least

OpenStudy (nincompoop):

go back to Pascal's triangle and there is actually a formula for it

OpenStudy (abbles):

Yes, I got 252 and I wanted to check to make sure it was right?

OpenStudy (abbles):

To determine the binomial coefficient for the fifth term in the expansion of (x + y)^10, I would start by finding the (n above r). To find the r value, which is the exponent/power x is being raised to in the fifth term, I subtracted 5 from 10 (the n value) to get 5. The n value is 10 because that is the exponent the binomial is being raised to. So the equation/whatever would be (10 above 5). The first step to solve that is to factor the factorials. 10!/5!5! = (10 x 9 x 8 x 7 x 6 x 5 x 4 x 3 x 2)/(5 x 4 x 3 x 2 x 5 x 4 x 3 x 2). I cancelled out some of the numbers until I got 252 over 1. Can you check my work @nincompoop

OpenStudy (abbles):

I'm homeschooled and working out of a textbook that doesn't show the answer for this. As I'm homeschooled, it isn't due so answer whenever you can. I'd like to know if I'm doing this right. Thanks!

OpenStudy (anonymous):

(x+y)^10 compare with (a+b)^n so.a=x,b=y,n=10,Now (r+1)^th=Tr+1=nCr,a^n-r,b^r where r+1=5,r=4,then put the above formula.

OpenStudy (abbles):

@tutulk is the answer 252? my work is above